# Gamma-ray burst collimated fireballs: long vs short bursts by angular momentum and mass ratio

### v2

@@ 5.5 Initial central density and length scale function in time
In homologous collapses, the baryon core centre density $\rho_{c}(t)$ increases as the length scale $a(t)$ (5.14) decreases. The temporal equation (5.17) governs how the scale length $a(t)$ monotonically decreases in time. At the beginning of homologous collapses, i.e., $t_{\rm initial}$, the initial length scales and centre density $\rho^{\rm in}_{c}\equiv\rho_{c}(t_{\rm initial})$ are undetermined, varying from one homologous collapsing stellar system to others. In solving differential equation (5.17), we chose the second constant of time integration by using boundary conditions at the end $t_{\rm final}$ of homologous collapses, which give the minimal length scale $a_{\rm min}$ and the time scale $\tau_{\rm min}$. We use the $a_{\rm min}$ “sound horizon” and $\tau_{\rm min}$ as the basic units for the temporal variables $a(t)$ and $t$. In such a way, the solution is written in a compact form,
@@
approaching their maximal values $\rho^{\rm max}_{c}=(3\gamma\kappa)^{-\frac{1}{\gamma-1}}$ and $1/\sqrt{3}$. The maximal centre density $\rho^{\rm max}_{c}=(3\gamma\kappa)^{-\frac{1}{\gamma-1}}$ is about $10$ times the nuclear saturation energy density $\rho_{0}\approx 2.4\times 10^{35}{\rm ergs}/{\rm cm}^{3}$, corresponding to the number density $n_{0}\approx 1.6\times 10^{38}/{\rm cm}^{3}$. The reason for choosing the maximal baryon core central density $\rho^{\rm max}_{c}\approx 10\rho_{0}$ in this article is the following. We consider a massive stellar baryon core at nuclear saturation density $\rho_{0}$, about the density of neutron stars. The baryon core conserves its total mass $M$ and continuously collapses to a compact object of size about $2GM$ of black hole horizon. In this case, its average density increases and becomes about $5\sim 8$ times the nuclear saturation energy density.
@@ 5.6 The rotation law of homologous collapse core
We are now in the position of postulating the rotation law for a rigidly and slowly rotating stellar core, which undergoes homologous gravitational collapse. Since the stellar rotating scale $\Omega^{-1}$ and collapsing timescales $\tau_{\rm grav}=R/\dot{R}$ are much larger than microscopic interaction and thermalization timescales, we approximately consider the gravitational collapse of rotating stellar core is an adiabatic process. Namely, at each collapse step at time $t$, the stellar core is approximately in an equilibrium state as if the collapsing stellar core adiabatically evolves in a time sequence of equilibrium states during gravitational collapse.
@@ 5.7 Total core mass and angular momentum conservations
which is constant in time. This means the initial values from the beginning of collapses determine $J/M$ values. The larger $J/M$ values are, the larger rotating angular velocities $\Omega_{c}(t)$ are, and the $a^{2}\Omega^{2}_{c}$ values are large at the end of homologous collapses. The ratio $J/M$ is a physically sensible parameter not only for massive rotating core collapses, but also for massive rotating binary coalescence. Therefore, we will use the ratio $J/M$ to characterize rotating systems that undergo homologous collapses.
@@ 5.8 Slowly rotating approximation
It must be emphasized that the theoretical framework is qualitatively valid only for $\Omega_{c}(t)J/M\ll 1$, since (i) we neglect the centrifugal potential $\phi_{c}$ in studying a homologous collapse in Secs. 5.4 and 5.5; (ii) we assume the rotation law (5.26) in homologous collapses. Nevertheless, we will see how the axial-symmetrically repulsive $\phi_{c}$ potential affects the spherical-symmetrically attractive potential $\phi$. The latter converts the gravitational binding energy into the photon-pair thermal energy of fireballs at the end of a homologous collapse [26]. We will then get an insight into the axial-symmetric distribution of photon-pair thermal energy (collimated fireball) when core rotations are present.
@@ 6.1 Spherical fireball temperature of photons and $e^{+}e^{-}$ pairs
In non-rotating cases centrifugal potential $\phi_{c}=0$, we obtain the baryon temperature $T$ by using the local virial theorem (3.5) and homologous solution (5.18). The photon-pair temperature is given by $T_{\gamma}\approx 0.21T$ (4.3),
@@
at photon-pair sphere centre $r=0$. We define the outer boundary $r_{\gamma}$ of the photon-pair sphere (spherical fireball) by $T(r_{\gamma})=2m_{e}$, the energy threshold of electron-positron pair productions. Compared $r_{\gamma}$ with the baryon core boundary $r_{s}$ in Fig. 1, we find $r_{\gamma}<r_{s}$, namely the photon-pair sphere (fireball) is formed inside the baryon core, see also Fig. 5.
A posterior, as a self-consistency check, we find that the baryon temperature $T$ and photon temperature $T_{\gamma}$ are much smaller than baryon mass ($T_{\gamma}\approx 0.21T\ll m$). The condition is filled for applying the virial theorem (3.3).
@@ 6.2 Collimated fireball temperature of photons and $e^{+}e^{-}$ pairs
In rotating cases centrifugal potential $\phi_{c}\not=0$ (5.26), we obtain the baryon temperature $T$ by using the local virial theorem (3.5) and homologous solution (5.18). The photon-pair temperature is given by $T_{\gamma}\approx 0.21T$ (4.3),
@@
The expression for $T_{\gamma}$ highlights the interplay between the gravitational and centrifugal potentials in determining the thermal energy distribution of the photon-pair plasma during the homologous collapse of the rotating stellar core. We describe the produced photon-pair plasma forms axial symmetric collimation in the cylindrical coordinates. For illustration, we use the parameters, we use parameters $(\gamma,\lambda_{m})=(1.23,8\times 10^{-5})$, the blue case in the list (5.20), and $a^{2}\Omega^{2}_{c}=0.05$ for a slowly rotating core.
@@
$r^{j}_{\gamma}(\theta)\leq r_{\gamma},\quad R^{j}_{\gamma}(\theta)=a(t)r^{j}_{\gamma}(\theta).$ (6.6)
The photon-pair sphere volume $V_{\gamma}$ is larger than the photon-pair collimated volume $V^{j}_{\gamma}$, see Fig. 5. In rotating cases, we plot the collimated boundary $r^{j}_{\gamma}(\theta)$ in Fig. 6. As a self-consistency check, we find that the baryon temperature $T$ and photon temperature $T_{\gamma}$ are much smaller than baryon mass ($T_{\gamma}\approx 0.21T\ll m$). It fills the condition for applying the virial theorem (3.3).
Figure 5: Following the colourmap of Fig. 4 for spherical and collimated fireballs, we compare the photon-pair fireball radius $r_{\gamma}$ and collimated fireball boundary $r^{j}_{\gamma}(\theta)$ (6.6), see Fig. 4, with the boundary $r_{s}\approx 36$ of the baryon core density $f^{3}(r)$ (5.12), see Fig. 1. The photon-pair fireball radius $r_{\gamma}\lesssim 30$ and $R_{\gamma}=a_{\rm min}r_{\gamma}\sim(10^{7}-10^{10})$ cm. The photon-pair collimated fireball boundary $r^{j}_{\gamma}(\theta)$ is given in Fig. 6. Note that the green zone and boundary indicate a baryon core, the blue (orange) zone and boundary indicate a spherical (collimated) fireball of photons and $e^{+}e^{-}$ pairs, and the boundary defines at the threshold $T_{\gamma}(r_{\gamma})=2m_{e}$ ($T_{\gamma}[r^{j}_{\gamma}(\theta)]=2m_{e}$) respectively. The baryon temperature zone and boundary follow the photon-pair temperature zone and boundary for $T_{\gamma}=0.21T$. These spatial distributions of fireballs and baryon core at the end of homologous collapses show $r_{s}>r_{\gamma}\geq r^{j}_{\gamma}(\theta)$. The collimated fireballs (orange zone) rotate along $\hat{z}$ axes with nontrivial angular momentum $J_{\gamma}$ (7.1). They act as initial configurations for ultra-relativistic hydrodynamic evolution (early launching and later propagation): Spherical expansion or axial-symmetric jet evolution outwards, proceeding internal and external shocks to produce GRBs.
Figure 6: The photon-pair collimated fireball boundary $r^{j}_{\gamma}(\theta)$ shows that (i) $\theta\rightarrow 0$ and $r^{j}_{\gamma}(\theta)$ approaches to the radius $r_{\gamma}\lesssim 30$ of the photon-pair fireball; (ii) $\theta\rightarrow\pi/2$ and $r_{\gamma}(\theta)$ approaches to $r^{j}_{\gamma}(\pi/2)\approx 1$ of the photon-pair collimated fireball (the cylindrical collimation) width $R^{j}_{\gamma}=a_{\rm min}r^{j}_{\gamma}(\pi/2)\sim(10^{6}-10^{8})$ cm, which is the maximal lateral size. The collimated fireball angle (7.2) $\theta_{\rm coll}\approx 1.8/30\approx 3.4^{\circ}$.
@@ 6.3 Collimated fireball energy density and size profiles
Both sizes $R_{\gamma}$ and $R^{j}_{\gamma}$ are much larger than the photon mean-free path $\lambda_{\gamma}=(\sigma_{\gamma}n_{\gamma})^{-1}$, namely $\sigma_{\gamma}n_{\gamma}R_{\gamma}\gg 1$ or $\sigma_{\gamma}n_{\gamma}R^{j}_{\gamma}\gg 1$. The photon-pair spherical and collimated fireballs are deeply opaque. On the other hand, we find in Fig. 4 that the photon-pair spherical and collimated fireballs’ temperature $T_{\gamma}$ can be well above the critical energy threshold $1.022m_{e}=$ 1.02 MeV of electron-positron ($e^{+}e^{-}$) pair production. The photon-pair spherical/collimated fireball energy densities $\rho_{\gamma}$ can be well above the critical energy density $\rho_{e}=m_{e}^{4}=5.93\times 10^{-11}\rho_{0}$ of electron-positron pairs. This means the photons are quickly thermalised to form an electron-position-photon plasma, as briefly discussed in Sec. 3.
Up to now, we have studied in this simplified model the processes that the gravitational binding energy is extracted and converted to a collimated and opaque fireball energy at the end of gravitational stellar collapses. The effect of an external magnetic field in such a process needs further investigation. These studies of extracting gravitational binding energy in stellar collapses contrast with the processes [22, 23, 24] of extracting rotational energy from spinning black holes.
@@ 6.4 Collimated fireball energy and spectrum peak
At the end of homologous collapse when the central sound velocity $v^{c}_{s}=\sqrt{1/3}$, we define integration
@@
In Eqs. (6.13) and (6.15), we express total photon-pair energy $E_{\gamma}$ and spectral peak $E_{p}$ in terms of the initial core centre density $\rho^{\rm in}_{c}=\rho_{c}(t_{\rm initial})$ at the time $t_{\rm initial}$ when the core initiates homologous collapse. The final core centre density is about 10 times the nuclear density $\rho^{\rm max}_{c}\approx 10\rho_{0}$. If the core mass $M$ and baryon number $N$ are fixed and conserved, the smaller initial core centre density $\rho^{\rm in}_{c}$ is, the more extended core mass density profile $\rho(r,t_{\rm initial})$ is, i.e., the larger $r_{s}$ is. As a consequence, more gravitational energy can be gained in collapse and converted to the photon-pair sphere/collimation energy $E_{\gamma}$ and number $N_{\gamma}$. In contrast, these quantities are also functions of the material binding energy parameter $\lambda$, the averaged thermal index $\gamma$ and rotation parameter ${\mathcal{A}}^{2}a^{2}\Omega_{c}^{2}\propto J/M$. The rotation represents the repulsive reactions against gravitational collapse and energy gain.
@@ 7.1 Angular momentum conservation making an ultra-relativistic jet
The hydrodynamical outflows of the collimated fireballs due to the huge photon-pair pressure $p_{\gamma}\sim\rho_{\gamma}$ are ultra-relativistic. The hydrodynamical outflows of the collimated fireballs due to the extremely high photon-pair pressure $p_{\gamma}\sim\rho_{\gamma}$ are ultra-relativistic. The collimated fireball outflow carries away the energy $E_{\gamma}$ and angular momentum $J_{\gamma}$ from collapsed systems, engulfing the outer material and forming an ultra-relativistic ejecta interacting with stellar matter, as described in the standard GRB scenario. The novelty is that the collimated outflows and ejecta launch ultra-relativistic jets of nontrivial angular momenta $J_{\gamma}$ of collimated fireballs. In later propagation, ultra-relativistic jets maintain collimated structure due to the conservation of their nontrivial angular momentum $J_{\gamma}\not=0$, analogously to a gyroscope effect. Numerical/simulation-based studies are necessary for verifying such dynamics. Analytical studies of ultra-relativistic jet launching and propagation and its stability in magnetic fields will be topics for separate articles. Notwithstanding, we present in the next section some qualitative discussions on the ultra-relativistic jet formation, early launching and later propagation.

### v3

@@ 5.5 Initial central density and length scale function in time
In homologous collapses, the baryon core centre density $\rho_{c}(t)$ increases as the length scale $a(t)$ (5.14) decreases. The temporal equation (5.17) governs how the scale length $a(t)$ monotonically decreases in time. At the beginning of homologous collapses, i.e., $t_{\rm initial}$, the initial length scales and centre density $\rho^{\rm in}_{c}\equiv\rho_{c}(t_{\rm initial})$ are undetermined, varying from one homologous collapsing stellar system to another. In solving differential equation (5.17), we chose the second constant of time integration by using boundary conditions at the end $t_{\rm final}$ of homologous collapses, which give the minimal length scale $a_{\rm min}$ and the time scale $\tau_{\rm min}$. We use the $a_{\rm min}$ “sound horizon” and $\tau_{\rm min}$ as the basic units for the temporal variables $a(t)$ and $t$. In such a way, the solution is written in a compact form,
@@
[…]
@@ 5.6 The rotation law of homologous collapse core
We are now in the position of postulating the rotation law for a rigidly and slowly rotating stellar core, which undergoes homologous gravitational collapse. Since the stellar rotating scale $\Omega^{-1}$ and collapsing timescales $\tau_{\rm grav}=R/\dot{R}$ are much larger than microscopic interaction and thermalization timescales, we approximately consider the gravitational collapse of the rotating stellar core to be an adiabatic process. Namely, at each collapse step at time $t$, the stellar core is approximately in an equilibrium state as if the collapsing stellar core adiabatically evolves in a time sequence of equilibrium states during gravitational collapse.
@@ 5.7 Total core mass and angular momentum conservations
which is constant in time. This means the initial values from the beginning of collapse determine $J/M$ values. The larger $J/M$ values are, the larger rotating angular velocities $\Omega_{c}(t)$ are, and the $a^{2}\Omega^{2}_{c}$ values are large at the end of homologous collapses. The ratio $J/M$ is a physically sensible parameter not only for massive rotating core collapses, but also for massive rotating binary coalescence. Therefore, we will use the ratio $J/M$ to characterize rotating systems that undergo homologous collapses.
@@ 5.8 Slowly rotating approximation
It must be emphasized that the theoretical framework is qualitatively valid only for $\Omega_{c}(t)J/M\ll 1$, since (i) we neglect the centrifugal potential $\phi_{c}$ in studying a homologous collapse in Secs. 5.4 and 5.5; (ii) we assume the rotation law (5.26) in homologous collapses. Nevertheless, we will see how the axial-symmetrically repulsive $\phi_{c}$ potential affects the spherically symmetrically attractive potential $\phi$. The latter converts the gravitational binding energy into the photon-pair thermal energy of fireballs at the end of a homologous collapse [36]. We will then get an insight into the axial-symmetric distribution of photon-pair thermal energy (collimated fireball) when core rotations are present.
We make a few more emphases to end this section. The analytically tractable rigid core assumption (5.8) is a poor approximation for collapsing systems with large angular momenta, particularly for fast rotating cores and binary neutron star mergers. The differential rotation law (angular-momentum distribution law), as a simple example (5.6) or other complex rotation laws, should play a more prominent role. In this case, the core rotation energy $U_{c}$ should dissipate to the internal energy $h$, the length scale function $a(t)$ Eqs. (5.26) and (6.5) should be altered. The axial symmetry along the rotating axis still holds and we may need two length scale functions $a(t)$ and $b(t)$ for an ellipsoid, because the part near the rotating axis with smaller centrifugal potentials more easily undergoes gravitational collapses than the part far from the rotating axis with larger centrifugal potentials. The angular momentum of gravitational collapsing systems gets redistributed, possibly forming a disc-like configuration, see reviews, for example, [57]. However, it has become a difficult issue to handle even with numerical simulations. Furthermore, ultra-relativistic jets along the direction of rotating axes carry away the angular momentum. This further favours an axially symmetric disc-like configuration with characteristic radiation patterns.
Moreover, we have not considered external magnetic field energy budgets [58] in gravitational collapsing systems, apart from the role of magnetic fields in the QED and QCD cascade multiplicity processes of photon production. Nevertheless, using the simplest model of rigid rotation, we obtain the total mass and angular-momentum conservation laws (5.7-5.30) and the essential quantity $J/M$ (5.30) that are valid for general systems, and qualitative effects and trends that shed some light on the long-standing problem and provide a physical insight into complex numerical simulations. Further investigations are needed in future work.
@@ 6.1 Spherical fireball temperature of photons and $e^{+}e^{-}$ pairs
In non-rotating cases, centrifugal potential $\phi_{c}=0$, we obtain the baryon temperature $T$ by using the local virial theorem (3.5) and homologous solution (5.18). The photon-pair temperature is given by $T_{\gamma}\approx 0.21T$ (4.3),
@@
at photon-pair sphere centre $r=0$. We define the outer boundary $r_{\gamma}$ of the photon-pair sphere (spherical fireball) by $T(r_{\gamma})=2m_{e}$, the energy threshold of electron-positron pair production. Compared $r_{\gamma}$ with the baryon core boundary $r_{s}$ in Fig. 1, we find $r_{\gamma}<r_{s}$, namely the photon-pair sphere (fireball) is formed inside the baryon core, see also Fig. 5.
A posterior, as a self-consistency check, we find that the baryon temperature $T$ and photon temperature $T_{\gamma}$ are much smaller than the baryon mass ($T_{\gamma}\approx 0.21T\ll m$). The condition is filled for applying the virial theorem (3.3).
@@ 6.2 Collimated fireball temperature of photons and $e^{+}e^{-}$ pairs
In rotating cases, centrifugal potential $\phi_{c}\not=0$ (5.26), we obtain the baryon temperature $T$ by using the local virial theorem (3.5) and homologous solution (5.18). The photon-pair temperature is given by $T_{\gamma}\approx 0.21T$ (4.3),
@@
The expression for $T_{\gamma}$ highlights the interplay between the gravitational and centrifugal potentials in determining the thermal energy distribution of the photon-pair plasma during the homologous collapse of the rotating stellar core. We describe the produced photon-pair plasma forms axial symmetric collimation in cylindrical coordinates. For illustration, we use the parameters $(\gamma,\lambda_{m})=(1.23,8\times 10^{-5})$, the blue case in the list (5.20), and $a^{2}\Omega^{2}_{c}=0.05$ for a slowly rotating core.
@@
$r^{j}_{\gamma}(\theta)\leq r_{\gamma},\quad R^{j}_{\gamma}(\theta)=a(t)r^{j}_{\gamma}(\theta),$ (6.6)
where the function $r^{j}_{\gamma}(\theta)$ is implicitly determined by
$\frac{2m_{e}}{m}$$\approx$ $\frac{0.21}{18}\Big{\{}\frac{1}{\gamma(\gamma-1)}\left[(7\gamma-6)f^{3(\gamma-1)}(r^{j}_{\gamma})-\frac{\lambda}{6}\gamma(r^{j}_{\gamma})^{2}\right]$ (6.7)
-- $\frac{1}{2}a^{2}\Omega^{2}_{c}(r^{j}_{\gamma})^{2}\sin^{2}\theta\Big{\}}.$
The photon-pair sphere volume $V_{\gamma}$ is larger than the photon-pair collimated volume $V^{j}_{\gamma}$, see Fig. 5. In rotating cases, we plot the collimated boundary $r^{j}_{\gamma}(\theta)$ in Fig. 6. As a self-consistency check, we find that the baryon temperature $T$ and photon temperature $T_{\gamma}$ are much smaller than the baryon mass ($T_{\gamma}\approx 0.21T\ll m$). It fills the condition for applying the virial theorem (3.3).
Figure 5: Following the colourmap of Fig. 4 for spherical and collimated fireballs, we compare the photon-pair fireball radius $r_{\gamma}$ and collimated fireball boundary $r^{j}_{\gamma}(\theta)$ (6.6), see Fig. 4, with the boundary $r_{s}\approx 36$ of the baryon core density $f^{3}(r)$ (5.12), see Fig. 1. The photon-pair fireball radius $r_{\gamma}\lesssim 30$ and $R_{\gamma}=a_{\rm min}r_{\gamma}\sim(10^{7}-10^{10})$ cm. The photon-pair collimated fireball boundary $r^{j}_{\gamma}(\theta)$ is given in Fig. 6. Note that the green zone and boundary indicate a baryon core, the blue (orange) zone and boundary indicate a spherical (collimated) fireball of photons and $e^{+}e^{-}$ pairs, and the boundary defines at the threshold $T_{\gamma}(r_{\gamma})=2m_{e}$ ($T_{\gamma}[r^{j}_{\gamma}(\theta)]=2m_{e}$) respectively. The baryon temperature zone and boundary follow the photon-pair temperature zone and boundary for $T_{\gamma}=0.21T$. These spatial distributions of fireballs and baryon core at the end of homologous collapses show $r_{s}>r_{\gamma}\geq r^{j}_{\gamma}(\theta)$. The collimated fireballs (orange zone) rotate along $\hat{z}$ axes with nontrivial angular momentum $J_{\gamma}$ (7.1). They act as initial configurations for ultra-relativistic hydrodynamic evolution (early launching and later propagation): Spherical expansion or axial-symmetric jet evolution outwards, proceeding with internal and external shocks to produce GRBs.
Figure 6: The photon-pair collimated fireball boundary $r^{j}_{\gamma}(\theta)$ shows that (i) $\theta\rightarrow 0$ and $r^{j}_{\gamma}(\theta)$ approaches to the radius $r_{\gamma}\lesssim 30$ of the photon-pair fireball; (ii) $\theta\rightarrow\pi/2$ and $r_{\gamma}(\theta)$ approaches $r^{j}_{\gamma}(\pi/2)\approx 1$ of the photon-pair collimated fireball (the cylindrical collimation) width $R^{j}_{\gamma}=a_{\rm min}r^{j}_{\gamma}(\pi/2)\sim(10^{6}-10^{8})$ cm, which is the maximal lateral size. The collimated fireball angle (7.2) $\theta_{\rm coll}\approx 1.8/30\approx 3.4^{\circ}$.
@@ 6.3 Collimated fireball energy density and size profiles
Both sizes $R_{\gamma}$ and $R^{j}_{\gamma}$ are much larger than the photon mean-free path $\lambda_{\gamma}=(\sigma_{\gamma}n_{\gamma})^{-1}$, namely $\sigma_{\gamma}n_{\gamma}R_{\gamma}\gg 1$ or $\sigma_{\gamma}n_{\gamma}R^{j}_{\gamma}\gg 1$. The photon-pair spherical and collimated fireballs are deeply opaque. On the other hand, we find in Fig. 4 that the photon-pair spherical and collimated fireballs’ temperature $T_{\gamma}$ can be well above the critical energy threshold $2m_{e}=1.02$ MeV of electron-positron ($e^{+}e^{-}$) pair production. The photon-pair spherical/collimated fireball energy densities $\rho_{\gamma}$ can be well above the critical energy density $\rho_{e}=m_{e}^{4}=5.93\times 10^{-11}\rho_{0}$ of electron-positron pairs. This means the photons are quickly thermalised to form an electron-position-photon plasma, as briefly discussed in Sec. 3.
Up to now, we have studied in this simplified model the processes which the gravitational binding energy is extracted and converted to a collimated and opaque fireball energy at the end of gravitational stellar collapses. The effect of an external magnetic field in such a process needs further investigation. These studies of extracting gravitational binding energy in stellar collapses contrast with the processes [32, 33, 34] of extracting rotational energy from spinning black holes.
@@ 6.4 Collimated fireball energy and spectrum peak
At the end of homologous collapse, when the central sound velocity $v^{c}_{s}=\sqrt{1/3}$, we define integration
@@
In Eqs. (6.14) and (6.16), we express total photon-pair energy $E_{\gamma}$ and spectral peak $E_{p}$ in terms of the initial core centre density $\rho^{\rm in}_{c}=\rho_{c}(t_{\rm initial})$ at the time $t_{\rm initial}$ when the core initiates homologous collapse. The final core centre density is about 10 times the nuclear density $\rho^{\rm max}_{c}\approx 10\rho_{0}$. If the core mass $M$ and baryon number $N$ are fixed and conserved, the smaller initial core centre density $\rho^{\rm in}_{c}$ is, the more extended core mass density profile $\rho(r,t_{\rm initial})$ is, i.e., the larger $r_{s}$ is. As a consequence, more gravitational energy can be gained in collapse and converted to the photon-pair sphere/collimation energy $E_{\gamma}$ and number $N_{\gamma}$. In contrast, these quantities are also functions of the material binding energy parameter $\lambda$, the averaged thermal index $\gamma$ and the rotation parameter ${\mathcal{A}}^{2}a^{2}\Omega_{c}^{2}\propto J/M$. The rotation represents the repulsive reactions against gravitational collapse and energy gain.
@@ 7.1 Angular momentum conservation making an ultra-relativistic jet
The hydrodynamical outflows of the collimated fireballs due to the huge photon-pair pressure $p_{\gamma}\sim\rho_{\gamma}$ are ultra-relativistic. The hydrodynamical outflows of the collimated fireballs due to the extremely high photon-pair pressure $p_{\gamma}\sim\rho_{\gamma}$ are ultra-relativistic. The collimated fireball outflow carries away the energy $E_{\gamma}$ and angular momentum $J_{\gamma}$ from collapsed systems, engulfing the outer material and forming an ultra-relativistic ejecta interacting with stellar matter, as described in the standard GRB scenario. The novelty is that the collimated outflows and ejecta launch ultra-relativistic jets of nontrivial angular momenta $J_{\gamma}$ of collimated fireballs. In later propagation, ultra-relativistic jets maintain collimated structure due to the conservation of their nontrivial angular momentum $J_{\gamma}\not=0$, analogously to a gyroscope effect. Numerical/simulation-based studies are necessary for verifying such dynamics. Analytical studies of ultra-relativistic jets launching and propagation, and their stability in magnetic fields, will be topics for separate articles. Notwithstanding, we present in the next section some qualitative discussions on the ultra-relativistic jet formation, early launching and later propagation.